me.angeli invested $5,000 at 8 and 1/2% interest for 9 months. how much interest did he earn on his investment? what is the total value of me.angeli’s investment after 9 months?
Answer by dkppathak(439)

mr.angeli invested $5,000 at 8 and 1/2% interest for 9 months. how much interest did he earn on his investment? what is the total value of mr.angeli’s investment after 9 months?
solution
time =9 month =9/12=3/4 years=0.75 year
interest =pxtxr/100
interest=5000x8.5x0.75/100=$318.75
total value after 9 months =5000+318.75=$5318.75
i